OpenStudy (anonymous):

Standard notation in a + bi; 4(cos(45) + isin(45)) what does it equal?

OpenStudy (anonymous):

a=4*cos(45) b=4*sin(45)

OpenStudy (anonymous):

cos(45)=sin(45)=1/sqrt(2)=sqrt(2)/2 so 2*sqrt(2) + 2*sqrt(2)i